# FUSE : A Ridge and Random Forest-Based Metric for Evaluating MT in Indigenous Languages

### v1

@@ 1 Introduction
MT has made significant advancements in recent years, largely driven by neural machine translation (NMT) models  [1]. However, evaluating the quality of translations remains a major challenge, particularly for low-resource Indigenous languages.Traditional MT evaluation metrics such as Bilingual Evaluation Understudy (BLEU)  [2] Translation Edit Rate (TER)  [3] , and Character n-gram F-score (ChrF)  [4] rely on surface-level token overlap, which fails to capture semantic correctness, fluency, and linguistic structure—critical factors in evaluating translations for morphologically rich and polysynthetic languages. Indigenous languages, such as Bribri, Guarani, and Nahuatl, exhibit unique linguistic characteristics that pose challenges for conventional MT evaluation  [5]. These languages often lack standardized orthography, leading to multiple valid translations  [6]. They feature lexical complexity, including polysynthesis and noun incorporation, which makes word segmentation and alignment with reference translations difficult  [7]. They also rely on phonetic variations, making strict token-level matching unreliable. Due to these factors, existing evaluation metrics struggle to provide reliable assessments of translation quality for Indigenous languages. While metrics such as BLEU and ChrF focus on exact token matches, they fail to account for phonetic and semantic similarities in morphologically rich languages.
Recent learning-based MT evaluation methods have demonstrated improved correlation with human judgments by incorporating semantic information from neural embeddings  [8],  [9]. However, these methods are not specifically designed for Indigenous languages, which require additional phonetic and structural considerations.
Our approaches integrate multiple linguistic and computational features, including lexical similarity using Levenshtein distance  [10], phonetic similarity using Metaphone  [11] and Soundex encoding  [12], semantic similarity using sentence embeddings from LaBSE  [13], and fuzzy token similarity to handle morphological variations  [14]. We train a linear regression model on human-annotated translation scores, optimizing feature weights to maximize alignment with human evaluation  [15]. Our results demonstrate that FUSE achieves higher Pearson and Spearman correlation  [16] with human evaluations compared to traditional MT metrics.
@@
Fig. 1: FUSE architecture combining linguistic features with hybrid regression for MT evaluation.
@@ 2.1 Rule-Based Metrics
Traditional rule-based evaluation metrics such as BLEU, TER, and ChrF  [17] rely on surface-level matching between candidate and reference translations. BLEU computes n-gram precision, but often fails to capture semantic adequacy or fluency, especially for morphologically rich languages  [18]. TER introduces edit-based alignment with support for word reordering but lacks deep linguistic modeling  [19]. ChrF improves robustness through character-level n-gram matching, making it better suited for languages with orthographic variation, though it still struggles with paraphrastic and semantic variation  [4].
@@ 2.2 Embedding-Based Metrics
Embedding-based metrics use contextual word or sentence representations to capture deeper semantic information. BLEURT  [20] fine-tunes pretrained BERT models on human-annotated MT quality data to produce sentence-level scores. COMET  [21] builds on multilingual transformers like XLM-R  [22] and incorporates both source and reference embeddings. TransQuest  [23] uses Siamese BERT  [24] networks to predict quality by comparing sentence pairs. These models outperform rule-based metrics in high-resource settings but remain data-hungry and often overlook features critical to low-resource or orthographically diverse languages.
@@ 2.3 Learning-Based Metrics
Learning-based metrics leverage supervised training on human-annotated translation quality data. Many of these metrics also incorporate contextual embeddings as input features. For example, COMET  [21] uses multilingual transformer embeddings (XLM-R)  [22]trained on direct assessment scores to predict translation quality. Similarly, BLEURT  [20] fine-tunes BERT for MT evaluation tasks, while TransQuest  [23] uses a Siamese architecture to model sentence-level similarity. These models achieve high correlation with human judgments in high-resource settings but often underperform in low-resource conditions due to their reliance on large training data and lack of sensitivity to phonetic or orthographic variation.
@@ 2.4 Quality Estimation (Reference-Free Metrics)
Quality Estimation (QE) aims to assess translation quality without relying on reference translations. Systems like QuEst++  [25] and recent neural QE models predict quality directly from source and hypothesis pairs. These models are especially useful in scenarios where references are unavailable or infeasible to generate. However, QE models also require substantial training data and have limited evaluation in the context of morphologically rich or under-resourced languages, such as those considered in this work  [26].
@@ 4.2 Approach 2: Feature-Enriched Similarity (DistilUSE)
In this approach, we compute a similarity score that integrates three core dimensions: lexical, phonetic, and semantic similarity. Lexical similarity is captured using normalized Damerau-Levenshtein distance, which quantifies surface-level edits between the reference ($r$) and hypothesis ($h$). Phonetic similarity is derived from Double Metaphone encodings  [27], comparing pronunciation-alike sequences. Finally, semantic similarity is computed using cosine similarity between sentence-level embeddings from the multilingual model  [28]. This method provides a more robust, language-agnostic similarity measure by incorporating both surface-level and deep semantic features. The final score is computed as a weighted sum of the three components:
@@ 4.4 Approach 4: Data-Driven Weighted Similarity via Regression
This approach employs a data-driven method to combine multiple similarity metrics by learning optimal weights through linear regression  [29]. For each pair of reference and hypothesis strings $(r,h)$, we extract four similarity features: lexical similarity $L(r,h)$ based on normalized Damerau–Levenshtein distance, phonetic similarity $P(r,h)$ computed using a combination of Metaphone and Soundex encodings, semantic similarity $S(r,h)$ derived from cosine similarity of LaBSE sentence embeddings  [30], and fuzzy token similarity $F(r,h)$ based on the normalized token sort ratio. These four features form the input vector $X(r,h)=[L(r,h),P(r,h),S(r,h),F(r,h)]$. Two separate linear regression models are trained using human-annotated semantic and fluency scores as targets. The first model learns weights $w_{\text{sem}}$ to predict semantic quality, while the second learns weights $w_{\text{flu}}$ for fluency. The final similarity score is computed by taking the average of the two predicted scores:
@@ 4.6 Approach 6: Ensemble Regression with Ridge and Gradient Boosting
In this approach, a hybrid ensemble method is employed by combining both linear and non-linear regression models to more accurately reflect human judgments of translation quality. For each reference–hypothesis pair $(r,h)$, a feature vector $X(r,h)=[L(r,h),\\ P(r,h),S(r,h),F(r,h)]$ is computed. Here, $L(r,h)$ represents the normalized Damerau–Levenshtein similarity, capturing character-level overlap. $P(r,h)$ denotes the phonetic similarity, derived from a combination of Metaphone and Soundex encodings. $S(r,h)$ measures semantic similarity using the cosine similarity between LaBSE sentence embeddings, while $F(r,h)$ is a fuzzy token similarity score based on the token sort ratio. All features are normalized using Min-Max scaling to ensure training stability.A Ridge regression model is trained to predict semantic scores, producing a weight vector $w_{\text{sem}}$. In parallel, a Gradient Boosting Regressor (GBR) models fluency scores non-linearly. The final score is computed as a weighted ensemble of the predictions, with 70% from the Ridge-based semantic score and 30% from the GBR-based fluency score. fluency score:
@@ 6 Results
On the development set, Approach 5 achieves the highest overall performance, attaining the best average Spearman (0.8001) and Pearson (0.8455) correlations across all three language pairs. This variant, visualized in Figure Fig. 1, employs a hybrid model combining Ridge regression (for semantic scoring) and Random Forest regression (for fluency), benefiting from the interpretability of linear models and the flexibility of ensemble-based non-linear modeling. Notably, it outperforms all other approaches on the Bribri language, likely due to its ability to capture intricate phonetic and lexical variability through feature learning. Approach 6, which replaces the fluency model with Gradient Boosting, performs comparably well—achieving top correlations for Guarani (Pearson: 0.8667) and Nahuatl (Spearman: 0.8216, Pearson: 0.8331)—suggesting that boosting methods are effective at modeling complex relationships in morphologically rich languages. In contrast, traditional feature-weighted approaches (Approaches 1–3) yield moderate results due to the absence of supervised weight optimization or the exclusive use of linear models, which limits their capacity to model non-linear dependencies. These development results are summarized in  Tab. 2.
On the held-out test set, a similar trend is observed: Approach 5 (RaaVa 2) achieves the highest average correlation with human annotations, ranking first in the shared task. As shown in  Tab. 3, this consistency across both development and test sets highlights the generalization capability of the ensemble architecture and validates the inclusion of phonetic, semantic, lexical, and fuzzy features. These findings further underscore the importance of integrating diverse linguistic signals with adaptive feature learning for MT evaluation in orthographically variable and low-resource Indigenous languages.
@@
  • Note: Best-performing scores in each column are highlighted in green and bold. Results are based on dev set correlations with human annotations.

### v2

@@ Abstract
FUSE : A Ridge and Random Forest-Based Metric for Evaluating MT in Indigenous Languages
Rahul Raja Carnegie Mellon University Stanford University LinkedInthanks: Work does not relate to position at LinkedIn.                        Arpita Vats Boston University Santa Clara University LinkedIn11footnotemark: 1
@@ 1 Introduction
MT has made significant advancements in recent years, largely driven by neural machine translation (NMT) models  Lyu et al. (2024). However, evaluating the quality of translations remains a major challenge, particularly for low-resource Indigenous languages.Traditional MT evaluation metrics such as Bilingual Evaluation Understudy (BLEU)  Papineni et al. (2002a) Translation Edit Rate (TER)  Snover et al. , and Character n-gram F-score (ChrF)  Popović rely on surface-level token overlap, which fails to capture semantic correctness, fluency, and linguistic structure—critical factors in evaluating translations for morphologically rich and polysynthetic languages. Indigenous languages, such as Bribri, Guarani, and Nahuatl, exhibit unique linguistic characteristics that pose challenges for conventional MT evaluation  Chen et al. (2023). These languages often lack standardized orthography, leading to multiple valid translations  Aepli et al. (2023). They feature lexical complexity, including polysynthesis and noun incorporation, which makes word segmentation and alignment with reference translations difficult  Tyers and Mishchenkova (2020). They also rely on phonetic variations, making strict token-level matching unreliable. Due to these factors, existing evaluation metrics struggle to provide reliable assessments of translation quality for Indigenous languages. While metrics such as BLEU and ChrF focus on exact token matches, they fail to account for phonetic and semantic similarities in morphologically rich languages.
Recent learning-based MT evaluation methods have demonstrated improved correlation with human judgments by incorporating semantic information from neural embeddings  Mathur et al. (2019),  Gumma et al. (2025). However, these methods are not specifically designed for Indigenous languages, which require additional phonetic and structural considerations.
Our approaches integrate multiple linguistic and computational features, including lexical similarity using Levenshtein distance  Levenshtein (1966), phonetic similarity using Metaphone  Philips (1990) and Soundex encoding  Russell (1918), semantic similarity using sentence embeddings from LaBSE  Feng et al. (2022), and fuzzy token similarity to handle morphological variations  Kondrak (2005). We train a linear regression model on human-annotated translation scores, optimizing feature weights to maximize alignment with human evaluation  Callison-Burch et al. (2006). Our results demonstrate that FUSE achieves higher Pearson and Spearman correlation  Spearman (1904) with human evaluations compared to traditional MT metrics.
@@
Figure 1: FUSE architecture combining linguistic features with hybrid regression for MT evaluation.
@@ 2.1 Rule-Based Metrics
Traditional rule-based evaluation metrics such as BLEU, TER, and ChrF  Popović (2015) rely on surface-level matching between candidate and reference translations. BLEU computes n-gram precision, but often fails to capture semantic adequacy or fluency, especially for morphologically rich languages  Papineni et al. (2002b). TER introduces edit-based alignment with support for word reordering but lacks deep linguistic modeling  Snover et al. (2006). ChrF improves robustness through character-level n-gram matching, making it better suited for languages with orthographic variation, though it still struggles with paraphrastic and semantic variation  Popović .
@@ 2.2 Embedding-Based Metrics
Embedding-based metrics use contextual word or sentence representations to capture deeper semantic information. BLEURT  Sellam et al. (2020) fine-tunes pretrained BERT models on human-annotated MT quality data to produce sentence-level scores. COMET  Rei et al. (2020) builds on multilingual transformers like XLM-R  Conneau et al. (2020) and incorporates both source and reference embeddings. TransQuest  Ranasinghe et al. (2020) uses Siamese BERT  Reimers and Gurevych (2019a) networks to predict quality by comparing sentence pairs. These models outperform rule-based metrics in high-resource settings but remain data-hungry and often overlook features critical to low-resource or orthographically diverse languages.
@@ 2.3 Learning-Based Metrics
Learning-based metrics leverage supervised training on human-annotated translation quality data. Many of these metrics also incorporate contextual embeddings as input features. For example, COMET  Rei et al. (2020) uses multilingual transformer embeddings (XLM-R)  Conneau et al. (2020)trained on direct assessment scores to predict translation quality. Similarly, BLEURT  Sellam et al. (2020) fine-tunes BERT for MT evaluation tasks, while TransQuest  Ranasinghe et al. (2020) uses a Siamese architecture to model sentence-level similarity. These models achieve high correlation with human judgments in high-resource settings but often underperform in low-resource conditions due to their reliance on large training data and lack of sensitivity to phonetic or orthographic variation.
@@ 2.4 Quality Estimation (Reference-Free Metrics)
Quality Estimation (QE) aims to assess translation quality without relying on reference translations. Systems like QuEst++  Specia et al. and recent neural QE models predict quality directly from source and hypothesis pairs. These models are especially useful in scenarios where references are unavailable or infeasible to generate. However, QE models also require substantial training data and have limited evaluation in the context of morphologically rich or under-resourced languages, such as those considered in this work  Sindhujan et al. (2025).
@@ 4.2 Approach 2: Feature-Enriched Similarity (DistilUSE)
In this approach, we compute a similarity score that integrates three core dimensions: lexical, phonetic, and semantic similarity. Lexical similarity is captured using normalized Damerau-Levenshtein distance, which quantifies surface-level edits between the reference ($r$) and hypothesis ($h$). Phonetic similarity is derived from Double Metaphone encodings  Yacob (2004), comparing pronunciation-alike sequences. Finally, semantic similarity is computed using cosine similarity between sentence-level embeddings from the multilingual model  Reimers and Gurevych (2019b). This method provides a more robust, language-agnostic similarity measure by incorporating both surface-level and deep semantic features. The final score is computed as a weighted sum of the three components:
@@ 4.4 Approach 4: Data-Driven Weighted Similarity via Regression
This approach employs a data-driven method to combine multiple similarity metrics by learning optimal weights through linear regression  Kuchibhotla et al. (2019). For each pair of reference and hypothesis strings $(r,h)$, we extract four similarity features: lexical similarity $L(r,h)$ based on normalized Damerau–Levenshtein distance, phonetic similarity $P(r,h)$ computed using a combination of Metaphone and Soundex encodings, semantic similarity $S(r,h)$ derived from cosine similarity of LaBSE sentence embeddings  Chimoto and Bassett (2022), and fuzzy token similarity $F(r,h)$ based on the normalized token sort ratio. These four features form the input vector $X(r,h)=[L(r,h),P(r,h),S(r,h),F(r,h)]$. Two separate linear regression models are trained using human-annotated semantic and fluency scores as targets. The first model learns weights $w_{\text{sem}}$ to predict semantic quality, while the second learns weights $w_{\text{flu}}$ for fluency. The final similarity score is computed by taking the average of the two predicted scores:
@@ 4.6 Approach 6: Ensemble Regression with Ridge and Gradient Boosting
In this approach, a hybrid ensemble method is employed by combining both linear and non-linear regression models to more accurately reflect human judgments of translation quality. For each reference–hypothesis pair $(r,h)$, a feature vector $X(r,h)=[L(r,h),P(r,h),S(r,h),F(r,h)]$ is computed. Here, $L(r,h)$ is the normalized Damerau–Levenshtein similarity capturing character-level overlap, $P(r,h)$ is the phonetic similarity derived from a combination of Metaphone and Soundex encodings, $S(r,h)$ is the cosine similarity between LaBSE sentence embeddings representing semantic similarity, and $F(r,h)$ is a fuzzy token similarity score based on the token sort ratio. All features are normalized using Min-Max scaling for training stability. A Ridge regression model is trained to predict semantic scores, producing a weight vector $w_{\text{sem}}$. In parallel, a Gradient Boosting Regressor (GBR) is trained to model fluency scores non-linearly. The final score is computed by taking a weighted ensemble of the predictions: 70% from the Ridge-based semantic score and 30% from the GBR-based fluency score:
@@ 6 Results
On the development set, Approach 5 achieves the highest overall performance, attaining the best average Spearman (0.8001) and Pearson (0.8455) correlations across all three language pairs. This variant, visualized in Figure fig. 1, employs a hybrid model combining Ridge regression (for semantic scoring) and Random Forest regression (for fluency), benefiting from the interpretability of linear models and the flexibility of ensemble-based non-linear modeling. Notably, it outperforms all other approaches on the Bribri language, likely due to its ability to capture intricate phonetic and lexical variability through feature learning. Approach 6, which replaces the fluency model with Gradient Boosting, performs comparably well—achieving top correlations for Guarani (Pearson: 0.8667) and Nahuatl (Spearman: 0.8216, Pearson: 0.8331)—suggesting that boosting methods are effective at modeling complex relationships in morphologically rich languages. In contrast, traditional feature-weighted approaches (Approaches 1–3) yield moderate results due to the absence of supervised weight optimization or the exclusive use of linear models, which limits their capacity to model non-linear dependencies. These development results are summarized in  Table 2.
On the held-out test set, a similar trend is observed: Approach 5 (RaaVa 2) achieves the highest average correlation with human annotations, ranking first in the shared task. As shown in  Table 3, this consistency across both development and test sets highlights the generalization capability of the ensemble architecture and validates the inclusion of phonetic, semantic, lexical, and fuzzy features. These findings further underscore the importance of integrating diverse linguistic signals with adaptive feature learning for MT evaluation in orthographically variable and low-resource Indigenous languages.
@@
  • Note: Best-performing scores in each column are highlighted in red and bold. Results are based on dev set correlations with human annotations.